OpenStudy (anonymous):

lim tanx as x approaches pi/2

Parth (parthkohli):

http://www.wolframalpha.com/input/?i=Tan+graph&dataset= Look at \(\pi \over 2\) in the graph.

OpenStudy (anonymous):

I believe the answer is 1 because in the graph there is asymptotes at -0.5pi, 0.5pi, 1.5pi, and 2.5pi and each asymptote is separated by 1 unit. But im not completely sure

Parth (parthkohli):

Nopee..

Parth (parthkohli):

I don't find it existing as a two-sided limit.

OpenStudy (anonymous):

well i believe they only want the answer from the positive direction

Parth (parthkohli):

From the right then.

OpenStudy (anonymous):

yes from the right

Parth (parthkohli):

If you plug in numbers like \[\tan\left({ \pi + 0.0000000000000000000000000001 \over 2}\right) \]You get a big, big number.

Parth (parthkohli):

Actually, a small, small one.

OpenStudy (anonymous):

Now why would you add that number to pi?

Parth (parthkohli):

Because you do NOT want tan(pi/2)... but tangent of something just a little bigger than pi/2... but still very small.

OpenStudy (anonymous):

So how would i use that to find the lim?